OpenStudy (anonymous):

Half Angle Formula help? Square ((1+cos 4x) / 2)

OpenStudy (anonymous):

cos 4x = 2cos^2(2x) - 1 so 1+ cos 4x = 1+2cos^2(2x) -1 = 2cos^2(2x)

OpenStudy (anonymous):

isnt is cos 2x?

OpenStudy (anonymous):

hmmm?

OpenStudy (anonymous):

cos 4x = cos 2(2x) right? so cos you can apply the formula of cos 2x for this

OpenStudy (anonymous):

oh you asked me for the final amswer??? yes it's cos 2x lol

OpenStudy (anonymous):

haha yes. thank you!

OpenStudy (anonymous):

oh I want to correct the answer again it should be the absolute of cos2x which is more correct loo